the fuel efficiency of an average car is given by y= -0.0177x*2 + 1.48x + 3.39 where x is the speed im miles per hour and y is the fuel economy in miles per gallon.
1. find the speed at which an average car will get 20 mi/gal
2. what speed yields the best fuel economy?
3. determine the range of speeds that yield 32 mi/gal or better


please answer
tHaNk yOU

Answer by stanbon(48535)

the fuel efficiency of an average car is given by y= -0.0177x*2 + 1.48x + 3.39 where x is the speed im miles per hour and y is the fuel economy in miles per gallon.
1. find the speed at which an average car will get 20 mi/gal
solve -0.0177x*2 + 1.48x + 3.39 = 20
---
If that equation is really a quadratic there are two solutions:
x = 70.26 mph and x = 13.36 mph
--------------------------------------------
2. what speed yields the best fuel economy?
x = -b/2a = -1.48/(2*-0.0177) = 41.81 mph
---------------------------------
3. determine the range of speeds that yield 32 mi/gal or better
33.34 <= x <= 53.28
=============================
Cheers,
Stan H.